mynamejeff13:

on my math homework i got the question 13 to the power of 7 times 32 can you help me

RhianaB112:

ok i am doing it rn

mynamejeff13:

thank you

RhianaB112:

np

RhianaB112:

1.1044277

mynamejeff13:

omg thank you, you saved my life

Cooltron:

hello, i have no idea how to do this.

mynamejeff13:

if i got that wrong i would have got an F

RhianaB112:

what?

RhianaB112:

hold on

Cooltron:

also that makes no sense

mynamejeff13:

i needed that answer to save my grade

Cooltron:

how could it be an answer less than 2?

RhianaB112:

ok i am working on it

Cooltron:

there is no division or subtraction in that math problem, so how could it be less than 2?

RhianaB112:

4*8 = 32 2*16 = 32 1*32 =32

mynamejeff13:

ill just ask google

RhianaB112:

1, 2, 4, 8, 16 and 32

RhianaB112:

do you get it?

Cooltron:

the answer is 9.6920035e+15